# UEFA Champions League entry is incentive incompatible again from the 2024/25 season

### v2

@@ Abstract
A tournament is called incentive incompatible if it allows for a situation where a team could be strictly better off by losing. This paper uncovers that the entry rules of the UEFA Champions League, the most prestigious club football competition in Europe, suffer from this shortcoming again from the 2024/25 season. Consequently, a match with misaligned incentives has been barely avoided in the 2023/24 German Bundesliga. Two straightforward solutions are recommended to safeguard the integrity of the competition. Fairness can be easily guaranteed by reversing the order of two articles in the UEFA Champions League regulation.
@@ 1 Introduction
Incentive compatibility or strategy-proofness—the two phrases are used as synonyms in the following—of a tournament means that each contestant should have appropriate incentives to perform under any circumstances. Unsurprisingly, incentive compatibility is among the most important tournament design criteria (Devriesere et al.,, 2024), and the objective for a sport organisation should be to eliminate any game that raises potential unfairness ex-ante (Scelles et al.,, 2024). At first glance, losing could never be a dominant strategy in sports. Indeed, if only one tournament is played, every reasonable ranking method satisfies strategy-proofness. However, the rules in the real world are usually more complex and occasionally create perverse incentives for the participants (Csató,, 2021).
According to the seminal paper of Dagaev and Sonin, (2018), the allocation of vacant slots can easily lead to incentive incompatibility. Indeed, this issue resulted in the violation of strategy-proofness for the UEFA Europea League entry until the 2014/15 season (Dagaev and Sonin,, 2018), for the UEFA Champions League entry between the 2015/16 and 2017/18 seasons (Csató,, 2019), and the UEFA Champions League seeding from the 2015/16 season (Csató, 2020b, ). The introduction of the UEFA Nations League has also allowed for games where a team has an incentive to lose (Csató,, 2021; Csató, 2022a, ; Haugen and Krumer,, 2021).
Here we show that, although the recent fundamental reform of the UEFA Champions League has restored the incentive compatibility of the seeding system, the misaligned order in which the vacancies are filled and the novel European performance spots are assigned leads to the violation of strategy-proofness. Furthermore, misaligned incentives almost occurred in the last round of the 2023/24 German Bundesliga in the match Eintracht Frankfurt vs. RB Leipzig. However, incentive compatibility can be guaranteed, for instance, by slightly modifying the UEFA Champions League regulation (UEFA,, 2024): the order of applying Articles 3.04 and 3.07 should be reversed. Therefore, UEFA is called to address the worrying deficiency as soon as possible.
The paper is structured as follows. Section 2 gives a concise summary of literature, focusing on the works that uncovered a situation where a team could have been strictly better off by losing. A motivating example is presented in Section 3. Section 4 discusses the root of the problem and offers potential solutions, while Section 5 concludes.
@@ 2 Related literature
Since tanking—the act of deliberately losing a game in order to gain some other advantage—is against the spirit of sports, incentive compatibility is an extensively discussed topic in the field of tournament design (Kendall and Lenten,, 2017; Csató,, 2021). Probably the most serious type of incentive incompatibility is when a team can be strictly better off by losing. According to our knowledge, such a possibility has been identified first by Dagaev and Sonin, (2018) in the UEFA Europa League. Several qualification tournaments for the FIFA World Cups and the UEFA European Championships have been demonstrated to suffer from the same weakness (Csató, 2020a, ), even allowing for a match where the dominant strategy of both opponents is to play a draw (Csató, 2020c, ).
In the past, UEFA club competitions violated strategy-proofness mainly due to the unfair allocation of vacant slots. Vacant slots occur if teams can qualify for a tournament in more than one way, thus, the organiser needs to specify what happens if at least two conditions of qualification are satisfied. An example is the UEFA Europa League: teams can enter via both the domestic championship and the domestic cup. According to Dagaev and Sonin, (2018), these tournament systems consisting of multiple round-robin and knockout tournaments are incentive incompatible if (1) there are at least two round-robin tournaments with at least one prize; (2) there is one round-robin tournament and at least one knockout tournament with at least one prize, and the vacant slots are not always filled through the round-robin tournament.
This theoretical finding might seem not especially interesting for sports administrators. However, the misaligned allocation policy has been responsible for the incentive incompatibility of
@@
Fortunately, Section 4 will reveal that the third problem has been solved by the new seeding rules effective from the 2024/25 season. On the other hand, the reform has also created potentially misaligned incentives regarding the UEFA Champions League entry as will be seen in Section 3.
Hence, our research is connected to the studies evaluating the effects of rule changes in the UEFA club competitions. Besides Csató, 2020b , the alternative seeding regime of the UEFA Champions League, introduced in the 2015/16 season, has been the subject of two further papers (Corona et al.,, 2019; Dagaev and Rudyak,, 2019). The modified structure of the champions path of the UEFA Champions League qualifying system has been examined by Csató, 2022b . A recent analysis of the away goals rule (Bahamonde-Birke and Bahamonde-Birke,, 2023) provides more insight into the consequences of abolishing this tie-breaking rule from the 2021/22 season. Last but not least, Gyimesi, (2024) uncovers how the new tournament format of the UEFA Champions League from the 2024/25 season affects competitive balance. The current paper is among the first ones that address this fundamental innovation in the most prestigious European club football competition.
@@ 3 A real-world example
Currently, the four highest-ranked associations—including Germany—are represented by four clubs in the UEFA Champions League. In addition, both the UEFA Champions League and UEFA Europa League titleholders have a guaranteed place in the league phase even if they do not qualify directly through their domestic championship. The possible vacancy is filled according to UEFA, (2024, Article 3.04):
@@
UEFA has introduced the so-called European performance spot from the 2024/25 season (UEFA,, 2024, Article 3.07):
@@ 7 SC Freiburg 11 9 13 44 56 --12 42
Table 1 shows the standing of the 2023/24 German Bundesliga on the morning of 18 May 2024 (the last matchday) with a slight modification: the result of the match Borussia Dortmund vs. 1899 Hoffenheim, played on 15 February 2024, is changed to 4-3 instead of the 2-3. At that moment, it was already known that Germany and Italy finished in the top two of the 2023/24 season association club coefficients, and earned the two European performance spots. The two clubs playing in the final of the UEFA Champions League on 1 June 2024 were known to be Borussia Dortmund and the Spanish club Real Madrid, too.
@@
  • If Borussia Dortmund is the fourth, then the first five teams qualify for the UEFA Champions League independently of the outcome of the 2023/24 UEFA Champions League final. If Dortmund wins the final, a vacancy is created, filled by the Ukrainian champion Shakhtar Donetsk (UEFA,, 2024, Article 3.04). If Dortmund loses the final, this team creates no vacancy.
@@
Although the above pathological situation has not been realised as Borussia Dortmund has been guaranteed to finish in the fifth place before the last round, UEFA has clearly had good luck in avoiding a strange game in the 2023/24 German Bundesliga. Some football fans have also noticed the problem of misaligned incentives for Eintracht Frankfurt, see https://m.sports.ru/tribuna/blogs/gazzzzzpacho/3239387.html.
4 Discussion
A recent literature survey on the causes and consequences of fraud in sport (Vanwersch et al.,, 2022) concludes that future research is needed to gather more knowledge on actionable elements with regards to the prevention of fraud. Our paper has examined the new format of the UEFA Champions League from this perspective. It should be acknowledged that the rules have improved from the 2024/25 season. Now the clubs are seeded based on their individual club coefficients except for automatically assigning the titleholder to the first pot (UEFA,, 2024, Article 16.01). Therefore, the regime used until the 2014/15 season has returned, and the seeding system satisfies incentive compatibility.
@@ 4 Discussion
UEFA, (2024, Article 3.04) has changed compared to the previous season of 2023/24. However, under both rules, no club could have an incentive to create a vacancy in order to obtain the released slot since either the highest-ranked not already qualified club from the association of the titleholder or a club from a different association benefits from the vacancy (see the cumbersome discussion of UEFA, (2024, Article 3.04b)). Consequently, Article 3.04 is not responsible for incentive incompatibility on its own.
On the other hand, the European performance spots are allocated only after the vacancies caused by the two titleholders are solved (UEFA,, 2024, Article 3.07). Thus, a team playing in the same round-robin tournament (domestic championship) as one of the (potential) titleholders can be interested in not creating vacancies in the UEFA Champions League. This can be achieved by “helping” other teams to be ranked over the (potential) titleholder—a clear example of a tanking opportunity as we have seen above. Such misaligned incentives can arise in the associations entitled to at least one spot in the league phase of the UEFA Champions League if they also earn one of the European performance spots. These are currently the 10 highest-ranked associations: in the 2024/25 season, England, Spain, Germany, Italy, France, the Netherlands, Portugal, Belgium, Scotland, and Austria.
@@
  • Rule change A: Modify the order of applying Articles 3.04 and 3.07 in the UEFA Champions League regulation (UEFA,, 2024). That is, first the two European performance spots are allocated (UEFA,, 2024, Article 3.07), and then the possible vacancies created by the titleholders are addressed (UEFA,, 2024, Article 3.04).
  • Rule change B: Modify the way of filling vacancies in UEFA, (2024, Article 3.04). That is, any vacancy created by a titleholder is filled by the highest-ranked not already qualified club from its association.
Both reforms satisfy incentive compatibility. For Rule change A, the argument is provided above. For Rule change B, strategy-proofness is implied by Dagaev and Sonin, (2018, Proposition 6) as all vacancies are filled through the only round-robin tournament, the domestic championship of the titleholder.
@@
If UEFA is not much worried about manipulation in a domestic league, and is not willing to change its regulations, the national associations are also able to mitigate the problem to some extent. Following the previous literature, the current paper has considered incentive incompatibility as a binary concept. However, the probability of a situation that is vulnerable to tanking can be quantified via Monte Carlo situations, analogously to Csató, 2022a , even if it would be difficult to implement.
Naturally, the set of matches to be played cannot be changed in a domestic championship without changing the tournament format, but the order of the matches is a decision variable. The scenario outlined in Section 3 has become threatening especially since the match Eintracht Frankfurt vs. RB Leipzig is played in the last round. The set of top teams can be relatively accurately estimated at the beginning of the season, hence, the last rounds can be scheduled to avoid any matches between the strongest teams. Such a schedule may lead to a tanking opportunity at the end of the season with a (much) lower probability. However, even if Eintracht Frankfurt had already played against RB Leipzig, it would be unfair if it turns out ex post that a team has missed qualifying for the UEFA Champions League due to winning a match months before. Therefore, it is the responsibility of UEFA to act.
@@ 5 Conclusions
Our paper has revealed that the UEFA Champions League rules, similar to three previous editions (Csató,, 2019), are incentive incompatible again from the 2024/25 season. The order in which the vacancies caused by the titleholders are filled and the two European performance spots are allocated has almost led to a game with misaligned incentives in the 2023/24 German Bundesliga. UEFA is strongly encouraged to adopt one of our recommendations in order to safeguard the integrity of its flagship tournament.

### v3

# Is UEFA really committed to sporting merit? A misaligned rule in the post-2024 Champions League

@@ Abstract
UEFA declares that it is fully committed to respecting the fundamental values of sport. However, the qualification of the post-2024 UEFA Champions League is not always based on sporting merit. Consequently, a match with misaligned incentives has been barely avoided in the 2023/24 German Bundesliga. Two straightforward solutions are recommended to safeguard the integrity of the competition. Fairness can be easily guaranteed by reversing the order of two articles in the Champions League regulation. Since UEFA repeatedly commits the same mistake that is widely known among researchers, governing bodies in sports are called to work more closely together with the scientific community.
@@ 1 Introduction
Ensuring the fairness and integrity of competitions is a key responsibility of sports policy (Sugden et al.,, 2023). Naturally, the Union of European Football Associations (UEFA) is also committed to these values, at least in declarations and media releases. For example, UEFA president Aleksander Čeferin has commented on the novel format of the UEFA Champions League with the following words (UEFA, 2024a, ):
UEFA has clearly shown that we are fully committed to respecting the fundamental values of sport and to defending the key principle of open competitions, with qualification based on sporting merit, fully in line with the values and solidarity-based European sports model.
However, is the UEFA Champions League qualification based on sporting merit under any circumstances? Is there any issue in its rules that compromises the fairness of this tournament and can potentially jeopardise competition?
The current note uncovers that, while UEFA may be committed to sporting merit in principles, their rules do not avoid an unfair game in a domestic league. In particular, although the recent fundamental reform of the UEFA Champions League has restored the integrity of the seeding system, the way in which the novel European performance spots are assigned allows for a situation when a team is strictly better off by losing. These misaligned incentives almost occurred in the last round of the 2023/24 German Bundesliga in the match Eintracht Frankfurt vs. RB Leipzig.
This problem is especially worrying because “the objective for a sport organisation should be to eliminate any game that raises potential unfairness ex-ante(Scelles et al.,, 2024, p. 535). Furthermore, its solution has been known since March 2013 when the first version of Dagaev and Sonin, (2018) become available.
@@ 2 Related literature
Since tanking—the act of deliberately losing a game—is against the spirit of sports, incentive compatibility is an extensively discussed topic in the field of tournament design (Kendall and Lenten,, 2017; Csató,, 2021). Probably the most serious type of incentive incompatibility is when a team can be strictly better off by losing. According to our knowledge, such a possibility has been identified first by Dagaev and Sonin, (2018) in the UEFA Europa League. Several qualification tournaments for the FIFA World Cups and the UEFA European Championships have been verified to suffer from the same weakness (Csató, 2020a, ), even allowing for a match where the dominant strategy of both opponents is to play a draw (Csató, 2020c, ).
In the past, UEFA club competitions violated strategy-proofness mainly due to the unfair allocation of vacant slots. Vacant slots occur if teams can qualify for a tournament in more than one way, thus, the organiser needs to specify what happens if at least two conditions of qualification are satisfied. A straightforward example is the UEFA Europa League, where teams can enter via both their domestic championship and their domestic cup. According to Dagaev and Sonin, (2018), these tournament systems consisting of multiple round-robin and knockout tournaments are incentive incompatible if (1) there are at least two round-robin tournaments with at least one prize; (2) there is one round-robin tournament and at least one knockout tournament with at least one prize, and the vacant slots are not always filled through the round-robin tournament.
This theoretical finding might seem not especially interesting for sports policy. However, the misaligned allocation rule has been responsible for the incentive incompatibility of
@@
Fortunately, Section 4 will reveal that the third problem has been solved by the new seeding rules effective from the 2024/25 season. On the other hand, the reform has also created potentially misaligned incentives in the UEFA Champions League qualification as will be seen in Section 3.
The introduction of the UEFA Nations League has also allowed for games where a team has an incentive to lose (Csató,, 2022; Haugen and Krumer,, 2021). Even though Scelles et al., (2024) identify no such games between 2018 and 2021, the authors call for more research around optimal competition formats that provide appropriate incentives for all teams in all games. Our paper is one of these follow-up studies.
@@ 3 A real-world example
Currently, the four highest-ranked associations—including Germany—are represented by four clubs in the UEFA Champions League. In addition, both the UEFA Champions League and UEFA Europa League titleholders automatically qualify for the league phase. The possible vacancy is filled according to UEFA, 2024b (, Article 3.04):
@@
UEFA has introduced the so-called European performance spot from the 2024/25 season (UEFA, 2024b, , Article 3.07):
@@ 7 SC Freiburg 11 9 13 44 56 --12 42
Table 1 shows the standing of the 2023/24 German Bundesliga on the morning of 18 May 2024 (the last matchday) with a slight modification: the result of the match Borussia Dortmund vs. 1899 Hoffenheim, played on 15 February 2024, is changed from 2-3 to 4-3. At that moment, it was already known that Germany and Italy finished in the top two of the 2023/24 season association club coefficients, and earned the two European performance spots. The two clubs playing in the final of the UEFA Champions League on 1 June 2024 were known to be Borussia Dortmund and the Spanish club Real Madrid, too.
@@
  • If Borussia Dortmund is the fourth, then the first five teams qualify for the UEFA Champions League independently of which team wins the 2023/24 UEFA Champions League final. If Dortmund wins the final, a vacancy is created, filled by the Ukrainian champion Shakhtar Donetsk (UEFA, 2024b, , Article 3.04). If Dortmund loses the final, this team creates no vacancy.
@@
The above pathological situation has not been realised as Borussia Dortmund has been guaranteed to finish in the fifth place before the last round. Nonetheless, UEFA has clearly had good luck in avoiding a strange game in the 2023/24 German Bundesliga. Some football fans have also noticed the problem of misaligned incentives for Eintracht Frankfurt, see https://m.sports.ru/tribuna/blogs/gazzzzzpacho/3239387.html.
4 Discussion and policy implications
A recent literature survey on fraud in sport (Vanwersch et al.,, 2022) concludes that future research is needed to accumulate more knowledge on the prevention of fraud. Our paper has examined the new format of the UEFA Champions League from this perspective. It should be acknowledged that the rules have improved from the 2024/25 season. Now the clubs are seeded based on their individual club coefficients except for the guaranteed place of the titleholder in the first pot (UEFA, 2024b, , Article 16.01). Therefore, the regime used until the 2014/15 season has returned, and the seeding system satisfies incentive compatibility.
@@ 4 Discussion and policy implications
UEFA, 2024b (, Article 3.04) has changed compared to the previous season of 2023/24. However, under both rules, no club could have an incentive to create a vacancy in order to obtain the released slot since either the highest-ranked not already qualified club from the association of the titleholder or a club from a different association benefits from the vacancy (see the cumbersome discussion of UEFA, 2024b (, Article 3.04b)). Consequently, Article 3.04 is not responsible for incentive incompatibility on its own.
On the other hand, the European performance spots are allocated only after the vacancies caused by the two titleholders are solved (UEFA, 2024b, , Article 3.07). Thus, a team playing in the same round-robin tournament (domestic championship) as one of the (potential) titleholders can be interested in not creating vacancies in the UEFA Champions League. This can be achieved by “helping” other teams to be ranked over the (potential) titleholder—a clear example of a tanking opportunity as we have seen above. Such misaligned incentives can arise in the associations entitled to at least one spot in the league phase of the UEFA Champions League if they also earn one of the European performance spots. These are currently the 10 highest-ranked associations: in the 2024/25 season, England, Spain, Germany, Italy, France, the Netherlands, Portugal, Belgium, Scotland, and Austria.
@@
  • Rule change A: Modify the order of applying Articles 3.04 and 3.07 in the UEFA Champions League regulation (UEFA, 2024b, ). That is, first the two European performance spots are allocated (UEFA, 2024b, , Article 3.07), and then the possible vacancies created by the titleholders are addressed (UEFA, 2024b, , Article 3.04).
  • Rule change B: Modify the way of filling vacancies in UEFA, 2024b (, Article 3.04). That is, any vacancy created by a titleholder is filled by the highest-ranked not already qualified club from its association.
Both reforms satisfy incentive compatibility. For Rule change A, the argument is provided above. For Rule change B, strategy-proofness is implied by Dagaev and Sonin, (2018, Proposition 6).
@@
If UEFA is not much worried about manipulation in a domestic league, and is not willing to change its regulations, the national associations are also able to mitigate the problem to some extent. Following the previous literature, the current paper has considered incentive incompatibility as a binary concept. However, the probability of a situation that is vulnerable to tanking can be quantified via Monte Carlo situations, analogously to Csató, (2022), even if it would be difficult to implement.
Naturally, the set of matches to be played cannot be changed in a domestic championship without changing the tournament format, but the order of the matches is a decision variable. The scenario outlined in Section 3 has become threatening especially since the match Eintracht Frankfurt vs. RB Leipzig is played in the last round. The set of top teams can be relatively accurately estimated at the beginning of the season, hence, the last rounds can be scheduled to avoid any matches between the strongest teams. Such a schedule may lead to a tanking opportunity at the end of the season with a (much) lower probability. However, even if Eintracht Frankfurt had already played against RB Leipzig, it would be unfair if it turns out ex post that a team has missed qualifying for the UEFA Champions League due to winning a match months before. Therefore, it is the responsibility of UEFA to act if it is really committed to its announced principles.
@@ 5 Conclusions
Our paper has revealed that the UEFA Champions League rules are incentive incompatible again from the 2024/25 season. The order in which the vacancies caused by the titleholders are filled and the two European performance spots are allocated has almost led to a game with misaligned incentives in the 2023/24 German Bundesliga. UEFA is strongly encouraged to adopt one of our recommendations in order to safeguard the integrity of its flagship tournament.
The issue presented above provides an instructive lesson for sports governing bodies around the world: commitment to sporting merit requires not only media releases but appropriately designed competitions. Since rule changes such as the introduction of the European performance spot can always affect fairness, a thorough assessment is necessary before they are implemented. In order to achieve this fundamental aim, Kendall and Lenten, (2017) suggest two alternative calls to action:
  • The governing bodies of major sports should invite academic representatives onto their committees, who would be tasked with identifying possible loopholes in proposed rule changes in consultation with the scientific community; or
  • All proposed rule changes should be posted in a public forum, where interested academics could comment, perhaps after running simulations.
Hopefully, UEFA and all other sports governing bodies will adopt one of these options in the future. This would be especially important because, as it has been highlighted in Section 2, UEFA repeatedly commits the same mistake in allocating vacant slots—while the fair solution has been provided by Dagaev and Sonin, (2018), and the first version of this paper is available since March 2013.